OpenStudy (anonymous):

Please help me PLEASE! Simplify the Radical expression. ^6 Squareroot 125x^3y^9z^6

OpenStudy (anonymous):

\[\sqrt[6]{27a ^{6}b ^{3}c ^{9}}\]

OpenStudy (anonymous):

I have to get this one right seriously

jigglypuff314 (jigglypuff314):

example -> ^6 squareroot(x^6) = x ^6 sqrt(x^12) = x^2 ^5 sqrt(x^5) = x ^5 sqrt(x^10) = x^(10/5) = x^2

OpenStudy (anonymous):

here is what i have but im pretty sure its not the answer \[27^{\frac{ 1 }{ 6 }}ab ^{\frac{ 1 }{ 2 }}c ^{\frac{ 3 }{ 2 }}\]

jigglypuff314 (jigglypuff314):

it's the right concept, but I would personally keep all the fraction stuff out of there like if you have ^6 sqrt(c^9) --> c* ^6sqrt(c^3)

jigglypuff314 (jigglypuff314):

so I only got a and c out of the sqrt the rest I just leave in there

OpenStudy (anonymous):

can you show me your answer you have no idea how bad i need to not get this wrong

jigglypuff314 (jigglypuff314):

I would prefer to not give answers but you've got the right idea! I was just suggesting to keep it to something like \[\huge \sqrt[6]{c^9} \rightarrow c~\sqrt[6]{c^3}\]

OpenStudy (anonymous):

so how about \[c \sqrt{bc}\]

jigglypuff314 (jigglypuff314):

erm for the ^6 sqrt (27 stuff) I got ac ^6sqrt(27b^3c^3)

OpenStudy (anonymous):

so like this? \[ac \sqrt[6]{27b ^{3}c ^{3}}\]

jigglypuff314 (jigglypuff314):

yes

OpenStudy (anonymous):

No thats not right

jigglypuff314 (jigglypuff314):

why not?

OpenStudy (anonymous):

because Im doing radical expressions it has to be simplified a different way, do you know of another way or rather can that equation be simplified more?

OpenStudy (anonymous):

I understand if your getting sick of this but you would be my hero if you helped me get this!

jigglypuff314 (jigglypuff314):

the only other way I can think of would be your fractional exponent stuff

OpenStudy (anonymous):

alright so if thats the case if my fractional exponents correct?

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

shoot it stated it as incorrect But you helped me alot and stayed with me to figure it out so heres a medal!

jigglypuff314 (jigglypuff314):

awww I'm sorry :/ I don't understand how it's wrong though :(

OpenStudy (anonymous):

these online courses are always weird with the numbers so ill just have to get the test unlocked tomorrow and re-do the tutorial again, but either way thank you very much!

jigglypuff314 (jigglypuff314):

oh well, good luck